help plzzExercise 5A-3 Cost Behavior; High-Low Method [LO5-10] Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singa
Answer #1

1.

Variable cost per unit = (174,000*$0.138) - (116,000*$0.177) / 174,000-116,000

Variable cost per unit = $24,012 - 20,532 / 58,000 = $0.06 per kilometer

Fixed cost = $24,012 - 10,440 (174,000*$0.06) = $13,572 per year

2.

Y = a + b X

Total cost = $13,572 + $0.06 * Number of kilometers driven

3.

Total operating cost = $13,572 + $0.06*145,000 = $22,272.